OpenStudy (rianiscool):

1:30 to reply will give a meld need help asap will give a medal need help asap will give a medal Jeff was counting the squirrels and birds in the park. He counted 16 heads and 44 legs. How many birds did Jeff count? A. 4 B. 8 C. 10 D. 11

OpenStudy (rianiscool):

is it 8

OpenStudy (rianiscool):

yes<>no

OpenStudy (rianiscool):

is it please

OpenStudy (anonymous):

Let b be the number of birds and s be the number of squirrels. We know that birds have two legs and squirrels have four. We therefore have the equations: 16 = b + s 44 = 2b + 4s From the first equation, we can derive that b = -s + 16. 1) 44 = 2(-s + 16) + 4s = -2s + 32 + 4s --> 2s = 12 --> s = 6 2) 16 = b + s = b + 6 --> b = 10 Check that our answer is true (b = 10, s = 6): 16 = 10 + 16 | TRUE 44 = 2(10) + 4(6) = 20 + 24 = 44 | TRUE There are ten birds, so the answer is (C).